# Spatially resolved H emission in B14-65666: compact starbursts, ionizing efficiency and gas kinematics in an advanced merger at the Epoch of Reionization

### v1

@@
11email: cprieto@cab.inta-csic.es 22institutetext: Departamento de FĂsica de la Tierra y AstrofĂsica, Facultad de Ciencias FĂsicas, Universidad Complutense de Madrid, E-28040, Madrid, Spain 33institutetext: Space Telescope Science Institute (STScI), 3700 San Martin Drive, Baltimore, MD 21218, USA 44institutetext: Department of Astronomy, Stockholm University, Oscar Klein Centre, AlbaNova University Centre, 106 91 Stockholm, Sweden 55institutetext: Centro de AstrobiologĂa (CAB), CSIC-INTA, Camino Bajo del Castillo s/n, E-28692 Villanueva de la Cañada, Madrid, Spain 66institutetext: Max Planck Institut fĂŒr Astronomie, Königstuhl 17, D-69117, Heidelberg, Germany 77institutetext: Kapteyn Astronomical Institute, University of Groningen, P.O. Box 800, 9700AV Groningen, The Netherlands 88institutetext: Cosmic Dawn Center (DAWN), Copenhagen, Denmark 99institutetext: Physikalisches Institut der UniversitĂ€t zu Köln, ZĂŒlpicher Str. 77, 50937 Köln, Germany 1010institutetext: European Space Agency, Space Telescope Science Institute, Baltimore, Maryland, USA 1111institutetext: DTU-Space, Elektrovej, Building 328, 2800, Kgs. Lyngby, Denmark 1212institutetext: DARK, Niels Bohr Institute, University of Copenhagen, 2200 Copenhagen, Denmark 1313institutetext: Telespazio UK for the European Space Agency (ESA), ESAC, Camino Bajo del Castillo s/n, 28692 Villanueva de la Cañada, Madrid, Spain 1414institutetext: European Space Agency (ESA), European Space Astronomy Centre (ESAC), Camino Bajo del Castillo s/n, 28692 Villanueva de la Cañada, Madrid, Spain 1515institutetext: School of Physics & Astronomy, Space Park Leicester, University of Leicester, 92 Corporation Road, Leicester LE4 5SP, UK 1616institutetext: Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721, USA 1717institutetext: Leiden Observatory, Leiden University, PO Box 9513, 2300 RA Leiden, The Netherlands 1818institutetext: Division of Physics, Faculty of Pure and Applied Sciences, University of Tsukuba,Tsukuba, Ibaraki 305-8571, Japan 1919institutetext: Tomonaga Center for the History of the Universe (TCHoU), Faculty of Pure and Applied Sciences, University of Tsukuba, Tsukuba, Ibaraki 305-8571, Japan 2020institutetext: National Astronomical Observatory of Japan, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan 2121institutetext: Waseda Research Institute for Science and Engineering, Faculty of Science and Engineering, Waseda University, 3-4-1 Okubo, Shinjuku, Tokyo 169-8555, Japan 2222institutetext: ETH ZĂŒrich, Institute for Particle Physics and Astrophysics, Wolfgang-Pauli-Str. 27, 8093 ZĂŒrich, Switzerland
@@
Spatially resolved H emission in B14-65666: compact starbursts, ionizing efficiency and gas kinematics in an advanced merger at the Epoch of Reionization
@@
We present MIRI/JWST medium resolution spectroscopy (MRS) and imaging (MIRIM) of B14-65666, a source identified as a Lyman-break and interacting galaxy at a redshift of â=â7.15. We detect the H line emission in this system, revealing a spatially-resolved structure of the H emitting gas, which consists of two distinct galaxies, E and W, at a projected distance of 0.4 arcsec apart (i.e. 2.2 kpc). One of the galaxies (E) is very compact (upper limit for the effective radius of 63 pc) in the rest-frame ultraviolet light, while the other galaxy (W) is more extended (effective radius of 348 pc), showing a clumpy structure reminiscent of a tidal tail. The total H luminosity implies that the system is forming stars at a rate of 76â±â8 Mââyr-1 and 30â±â4 Mââyr-1 for E and W galaxies, respectively. The ionizing photon production efficiency, ), for galaxies E and W has values of 25.1â±â0.1 Hz erg-1 and 25.5â±â0.1 Hz erg-1, within the range measured in galaxies at similar redshifts. The high values derived for the H equivalent widths (832â±â100 and 536â±â78 ) and the distinct locations of the E and W galaxies in the ) equivalent width (H) plane, indicate that the system is dominated by a young (less than 10 Myr) stellar population. The overall spectral energy distribution suggests that in addition to a young stellar population, the two galaxies may have mature (more than 100 Myr) stellar population and very different dust attenuation with galaxy E showing a larger attenuation (AVâ=â1.5 mag), compared to the almost dust-free (AVâ=â0.1 mag) galaxy W. The derived Star Formation Rate (SFR) and stellar masses identify the two galaxies as going through a starburst phase characterized by a specific SFR (sSFR) of 40âââ50âGyr-1. Galaxy E has an extreme stellar mass surface density (6ââ104âMââpc-2), close to that of the nuclei of low- galaxies while galaxy W (103âMâ pc-2) is consistent with the surface densities measured in galaxies at these redshifts. The kinematics of the ionized gas traced by the H line show a velocity difference of 175ââ28 kmâs-1 between the two components of B14-65666, and a broader profile for galaxy W (312â44 kmâs-1) relative to galaxy E (243ââ41 kmâs-1). The detailed study of B14-65666 shows the complex stellar and Interstellar Medium structure in merging galaxy systems was already in place by the Epoch of Reionization. The general properties of B14-65666 agree with those predicted for massive merging systems at redshifts of 7 and above in the FIRSTLIGHT cosmological simulations. The in-depth study of systems like B14-65666 reveal how galaxy mergers in the early Universe drive intense star formation, shape the interstellar medium, and influence the buildup of stellar mass, just 700âââ800 Myr after the Big Bang.
@@
The James Webb Space Telescope (JWST, Gardner et al. 2023), is transforming our understanding of galaxy formation, particularly during the so-called Epoch of Reionization (EoR). JWST has photometrically detected potential galaxy candidates up to redshift 25 (PĂ©rez-GonzĂĄlez et al. 2025), and spectroscopically confirmed galaxies up to redshift 14.44 (Naidu et al. 2025). At these redshifts, the rest-frame optical emission lines and continuum are redshifted into the mid-infrared (5âââ25âm), the spectral range covered by the Mid-Infrared Instrument (MIRI, Rieke et al. 2015; Wright et al. 2015, 2023). At redshifts of 7 and above, this instrument allows us to explore the strong optical emission lines, such as H.
@@
Mergers are expected to have a greater impact on galaxy evolution at higher redshifts than at lower redshifts, as the merger rate increases with redshift (Duan et al. 2025). The gravitational forces from mergers can cause significant disruptions in galaxy structures and movements, leading to features such as tidal tails and asymmetric stellar envelopes (e.g.Toomre & Toomre 1972; Mihos & Hernquist 1996; Bushouse et al. 2002; Conselice et al. 2003; Spilker et al. 2022). These disturbances can cause gas to flow from the outer regions into the center of galaxies, triggering the formation of new stars in central areas (PĂ©rez-GonzĂĄlez et al. 2024c). Simulations such as FirstLight (Ceverino et al. 2017) show that about 10% of galaxies at redshifts 6âââ10 exhibit clumpy structures, primarily formed through mergers. The central clumps are primarily composed of older stellar populations exceeding 50âMyr, while off-center clumps originate from gas debris along tidal tails during major mergers. These mergers have a short life, typically merging into the main galaxy within a few tens of million years (Nakazato et al. 2024). Observational studies also support the presence of mergers in the EoR, with recent works identifying clear merger signatures in high-redshift galaxies (e.g. Ălvarez-MĂĄrquez et al. 2023; Arribas et al. 2024; Harshan et al. 2024; Hu et al. 2025).
@@
High- star-forming galaxies (SFGs) are characterized by optical lines with high (rest-frame) values of the equivalent widths (EW up to 3000â) in extreme emission line galaxies (Boyett et al. 2024). There is also a trend of increasing EW with redshift which indicates an evolution in the galaxy specific star formation rate (Rinaldi et al. 2023; Rinaldi et al. 2025) and young stellar populations (ÂĄâ10âMyr). Studies at low redshifts have shown that galaxies in close pairs exhibit an SFR increase compared to isolated galaxies, highlighting the impact of gravitational interactions on gas dynamics and star formation (Barton et al. 2000). This enhancement is also observed at higher redshifts (6.5âââ8.5), with an excess in SFR of 0.26â±â0.11 dex found in very close pairs, with no significant increase at larger separations (Duan et al. 2024; PuskĂĄs et al. 2025).
@@
B14-65666 (also known as the galaxy âBig Three Dragonsâ) was discovered based on wide-field imaging data of the UltraVISTA survey of the Cosmological Evolution Survey (COSMOS) field in Y, J, H and Ks near-infrared bands (Bowler et al. 2012, Bowler et al. 2014). It was then spectroscopically detected in Ly (Furusawa et al. 2016). An image captured with the Hubble Space Telescope (HST) showed that the galaxy consists of two clearly separated components in the rest-frame ultraviolet (UV), thus suggesting that the object could be undergoing a merger of two galaxies (Bowler et al. 2017). Observations performed with Atacama Large Millimeter/submillimeter Array (ALMA) with Band 6 reported the detection of a â160âm dust continuum at the position of the components (Bowler et al. 2018). It was also detected [OâIII]â88âm, [CâII]â158âm, and underlying dust continuum emission at â90âm, in Bands 6 and 8 (Hashimoto et al. 2019). Furthermore, they showed that [OâIII]â88âm and [CâII]â158âm can be decomposed into two components spatially resolved and kinematically separated. These components are associated with the two UV emission peaks identified in the HST images. Additional observations with ALMA Band 7 did not detect the presence of the far-infrared [NâII]â122âm fine-structure emission line, nor did ALMA Band 3 observations detect other lines such as CO (6-5), CO (7-6), and [CâI] (2-1) (Sugahara et al. 2021; Hashimoto et al. 2023).
@@
B14-65666 is one of the few galaxies identified as undergoing a major merger during the Epoch of Reionization. This system allows the investigation of the physical properties of an early major merger, including its stellar structure, as well as the different phases of the Interstellar Medium (ISM). Additionally, this system provides a reference for direct comparisons with simulations. In this study, we present the first spatially resolved spectroscopy of the H emission in the system B14-65666 at â=â7.15 based on new MIRI/JWST integral field spectroscopy. The paper presents several of the physical properties of B14-65666 based on the combination of the new MIRI data with ancillary NIRCam imaging and ALMA spectroscopy. This paper is organized as follows. In Sect. 2 we present the observations and data calibrations. In Sect. 3 we outline our analysis methods of spectroscopy and photometry. In Sect. 4 we present the results of the analysis. This section includes the subsections in which we study the Spectral Energy Distribution of the galaxy (4.1), the attenuation (4.2), the star formation rate and burstiness (4.3), the H EW and ionizing photon production efficiency (4.4), the kinematics (4.5), stellar and dynamical mass (4.6), the mass-size relation (4.7), the starburst nature (4.8), and finally we analyze the origin and evolution of the system according to simulations (4.9). We discuss and analyze the spectral and spatial distribution of the H line and the derived properties of the galaxies and we study the morphology and SFR of the galaxy and compare it with local galaxies. Our conclusions are presented in Sect. 5.
In this paper we assume a flat CDM cosmology with â=â0.310, and â=â67.7âkmâs-1âMpc-1 (Planck Collaboration et al., 2020). With these parameters, 1ââ is equal to 5.275 kpc at â=â7.15 and the luminosity distance is =â72276.9 Mpc. For these cosmological parameters and redshift, the age of the Universe corresponds to 740 Myr. We use vacuum emission line wavelengths throughout the paper.
@@ 2.1 MIRI
B14-65666 was observed with MIRI on the 30th of April of 2023 as a part of the MIRI European Consortium guaranteed time observations (program ID 1284, PI: L. Colina). The observations are composed of MIRI imaging (MIRIM, Bouchet et al. 2015; Dicken et al. 2024) and integral field spectroscopy using the Medium Resolution Spectrograph (MRS, Wells et al. 2015; Argyriou et al. 2023).
@@ 2.1.1 MIRI spectroscopy
The MRS observations were performed with the SHORT band for channels 1âââ4. We only use Channel 1 in this study. This channel covers the wavelengths 4.9âââ5.74âm, which targets the H emission line at redshift 7.15, and other fainter emission lines, such as [NâII]â6550,6585â and [SâII]â6718,6733â. The total on-source exposure time of the MRS observations is 22743 seconds. The positive and negative 4-point dither patterns for point sources are used. For each dither position, a total of 6 integrations are obtained using the SLOWR1 readout mode with 19 groups each. When the MRS is targeting B14-65666, a parallel field is imaged using MIRI F770W filter. This imaging data is utilized exclusively for the astrometric calibration of the MRS (see Sect. 2.4).
We use version 1.15.1 of the JWST calibration pipeline and context 1276 of the Calibration Reference Data System (CRDS) to process and calibrate the MRS observations of B14-65666. The standard MRS pipeline procedure (Bushouse et al. 2024) is used with additional customized steps to improve the quality of the final MRS calibrated products and background (see Ălvarez-MĂĄrquez et al. 2023, 2024 for details). Additionally, we implement the algorithm developed by Spilker et al. (2023) to correct the striping pattern produced by the cosmic ray showers in the spectral cube. The final Channel 1 cube has a spatial and spectral sampling of 0.13âââ0.13âââ0.8ânm (Law et al. 2023), and the resolving power () is about 3500 (Labiano et al. 2021; Jones et al. 2023).

### v2

# Spatially resolved H emission in B14-65666: Compact starbursts, ionizing efficiency, and gas kinematics in an advanced merger at the Epoch of Reionization

@@
11email: cprieto@cab.inta-csic.es 22institutetext: Departamento de FĂsica de la Tierra y AstrofĂsica, Facultad de Ciencias FĂsicas, Universidad Complutense de Madrid, E-28040, Madrid, Spain 33institutetext: Space Telescope Science Institute (STScI), 3700 San Martin Drive, Baltimore, MD 21218, USA 44institutetext: Department of Astronomy, Stockholm University, Oscar Klein Centre, AlbaNova University Centre, 106 91 Stockholm, Sweden 55institutetext: Centro de AstrobiologĂa (CAB), CSIC-INTA, Camino Bajo del Castillo s/n, E-28692 Villanueva de la Cañada, Madrid, Spain 66institutetext: Leiden Observatory, Leiden University, PO Box 9513, 2300 RA Leiden, The Netherlands 77institutetext: Kapteyn Astronomical Institute, University of Groningen, P.O. Box 800, 9700AV Groningen, The Netherlands 88institutetext: Cosmic Dawn Center (DAWN), Copenhagen, Denmark 99institutetext: Physikalisches Institut der UniversitĂ€t zu Köln, ZĂŒlpicher Str. 77, 50937 Köln, Germany 1010institutetext: European Space Agency, Space Telescope Science Institute, Baltimore, Maryland, USA 1111institutetext: DTU-Space, Elektrovej, Building 328, 2800, Kgs. Lyngby, Denmark 1212institutetext: DARK, Niels Bohr Institute, University of Copenhagen, 2200 Copenhagen, Denmark 1313institutetext: Telespazio UK for the European Space Agency (ESA), ESAC, Camino Bajo del Castillo s/n, 28692 Villanueva de la Cañada, Madrid, Spain 1414institutetext: European Space Agency (ESA), European Space Astronomy Centre (ESAC), Camino Bajo del Castillo s/n, 28692 Villanueva de la Cañada, Madrid, Spain 1515institutetext: School of Physics & Astronomy, Space Park Leicester, University of Leicester, 92 Corporation Road, Leicester LE4 5SP, UK 1616institutetext: Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721, USA 1717institutetext: Max Planck Institut fĂŒr Astronomie, Königstuhl 17, D-69117, Heidelberg, Germany 1818institutetext: Division of Physics, Faculty of Pure and Applied Sciences, University of Tsukuba,Tsukuba, Ibaraki 305-8571, Japan 1919institutetext: Tomonaga Center for the History of the Universe (TCHoU), Faculty of Pure and Applied Sciences, University of Tsukuba, Tsukuba, Ibaraki 305-8571, Japan 2020institutetext: National Astronomical Observatory of Japan, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan 2121institutetext: Waseda Research Institute for Science and Engineering, Faculty of Science and Engineering, Waseda University, 3-4-1 Okubo, Shinjuku, Tokyo 169-8555, Japan 2222institutetext: ETH ZĂŒrich, Institute for Particle Physics and Astrophysics, Wolfgang-Pauli-Str. 27, 8093 ZĂŒrich, Switzerland
@@
Spatially resolved H emission in B14-65666: Compact starbursts, ionizing efficiency, and gas kinematics in an advanced merger at the Epoch of Reionization
@@
We present MIRI/JWST medium-resolution spectroscopy (MRS) and imaging (MIRIM) of B14-65666, a source identified as a Lyman-break and interacting galaxy at a redshift of â=â7.15. We detect the H line emission in this system, revealing a spatially resolved structure of the H-emitting gas, which consists of two distinct galaxies, E and W, at a projected distance of 0.4 arcsec apart (i.e., 2.2 kpc). One of the galaxies (E) is very compact (upper limit for the effective radius of 63 pc) in the rest-frame ultraviolet light, while the other galaxy (W) is more extended (effective radius of 348 pc), showing a clumpy structure reminiscent of a tidal tail. The total H luminosity implies that the system is forming stars at a rate of 76â±â8 Mââyr-1 and 30â±â4 Mââyr-1 for E and W galaxies, respectively. The ionizing photon production efficiency, ), for galaxies E and W, has values of 25.1â±â0.1 Hz erg-1 and 25.5â±â0.1 Hz erg-1, which is within the range measured in galaxies at similar redshifts. The high values derived for the H equivalent widths (832â±â100 and 536â±â78 ) and the distinct locations of the E and W galaxies in the ) equivalent width (H) plane indicate that the system is dominated by a young (under 10 Myr) stellar population. The overall spectral-energy distribution suggests that in addition to a young stellar population, the two galaxies may have mature (over 100 Myr) stellar populations and very different dust attenuations, with galaxy E showing a larger attenuation (AVâ=â1.5 mag) compared to the almost dust-free (AVâ=â0.1 mag) galaxy W. The derived star formation rate (SFR) and stellar masses identify the two galaxies as going through a starburst phase characterized by a specific SFR (sSFR) of 40âââ50âGyr-1. Galaxy E has an extreme stellar mass surface density (6ââ104âMââpc-2), close to that of the nuclei of low- galaxies, while galaxy W (103âMâ pc-2) is consistent with the surface densities measured in galaxies at these redshifts. The kinematics of the ionized gas traced by the H line show a velocity difference of 175ââ28 kmâs-1 between the two components of B14-65666 and a broader profile for galaxy W (312â44 kmâs-1) relative to galaxy E (243ââ41 kmâs-1). The detailed study of B14-65666 shows that the complex stellar and interstellar medium structure in merging galaxy systems was already in place by the Epoch of Reionization. The general properties of B14-65666 agree with those predicted for massive merging systems at redshifts of 7 and above in the FIRSTLIGHT cosmological simulations. The in-depth study of systems such as B14-65666 reveal how galaxy mergers in the early Universe drive intense star formation, shape the interstellar medium, and influence the buildup of stellar mass, just 700âââ800 Myr after the Big Bang.
@@
The James Webb Space Telescope (JWST, Gardner et al. 2023) is transforming our understanding of galaxy formation, particularly during the so-called Epoch of Reionization (EoR). The JWST has photometrically detected potential galaxy candidates up to redshift 25 (PĂ©rez-GonzĂĄlez et al. 2025) and spectroscopically confirmed galaxies up to redshift 14.44 (Naidu et al. 2025). At these redshifts, the rest-frame optical emission lines and continuum are redshifted into the mid-infrared (5âââ25âm), which is the spectral range covered by the Mid-Infrared Instrument (MIRI, Rieke et al. 2015; Wright et al. 2015, 2023). At redshifts of 7 and above, this instrument allows us to explore the strong optical emission lines, such as H.
@@
Mergers are expected to have a greater impact on galaxy evolution at higher redshifts than at lower redshifts, as the merger rate increases with redshift (Duan et al. 2025). The gravitational forces from mergers can cause significant disruptions in galaxy structures and movements, leading to features such as tidal tails and asymmetric stellar envelopes (e.g., Toomre & Toomre 1972; Mihos & Hernquist 1996; Bushouse et al. 2002; Conselice et al. 2003; Spilker et al. 2022). These disturbances can cause gas to flow from the outer regions into the center of galaxies, triggering the formation of new stars in central areas (PĂ©rez-GonzĂĄlez et al. 2024c). Simulations such as FirstLight (Ceverino et al. 2017) show that about 10% of galaxies at redshifts of 6âââ10 exhibit clumpy structures, which are primarily formed through mergers. The central clumps are mainly composed of older stellar populations exceeding 50âMyr, while off-center clumps originate from gas debris along tidal tails during major mergers. These mergers have a short life, typically merging into the main galaxy within a few tens of million years (Nakazato et al. 2024). Observational studies also support the presence of mergers in the EoR, with recent works identifying clear merger signatures in high-redshift galaxies (e.g., Ălvarez-MĂĄrquez et al. 2023; Arribas et al. 2024; Harshan et al. 2024; Hu et al. 2025).
@@
High- star-forming galaxies (SFGs) are characterized by optical lines with high (rest-frame) values of the equivalent widths (EW up to 3000â) in extreme emission line galaxies (Boyett et al. 2024). There is also a trend of EW increasing with redshift, which indicates an evolution in the galaxy-specific star formation rate (Rinaldi et al. 2023; Rinaldi et al. 2025) and young stellar populations (ÂĄâ10âMyr). Studies at low redshifts have shown that galaxies in close pairs exhibit an SFR increase compared to isolated galaxies, highlighting the impact of gravitational interactions on gas dynamics and star formation (Barton et al. 2000). This enhancement is also observed at higher redshifts (6.5âââ8.5), with an excess in SFR of 0.26â±â0.11 dex found in very close pairs, with no significant increase at larger separations (Duan et al. 2024; PuskĂĄs et al. 2025).
@@
B14-65666 (also known as the Big Three Dragons galaxy) was discovered based on wide-field imaging data of the UltraVISTA survey of the Cosmological Evolution Survey (COSMOS) field in Y, J, H, and Ks near-infrared bands (Bowler et al. 2012, Bowler et al. 2014). It was then spectroscopically detected in Ly (Furusawa et al. 2016). An image captured with the Hubble Space Telescope (HST) showed that the galaxy consists of two clearly separated components in the rest-frame ultraviolet (UV), thus suggesting that the object could be undergoing a merger of two galaxies (Bowler et al. 2017). Observations performed with the Atacama Large Millimeter/submillimeter Array (ALMA) with Band 6 reported the detection of an â160âm dust continuum at the position of the components (Bowler et al. 2018). It also detected [OâIII]â88âm, [CâII]â158âm, and underlying dust continuum emission at â90âm in Bands 6 and 8 (Hashimoto et al. 2019). Furthermore, they showed that [OâIII]â88âm and [CâII]â158âm can be decomposed into two components that are spatially resolved and kinematically separated. These components are associated with the two UV emission peaks identified in the HST images. Additional observations with ALMA Band 7 did not detect the presence of the far-infrared [NâII]â122âm fine-structure emission line, nor did ALMA Band 3 observations detect other lines such as CO (6-5), CO (7-6), and [CâI] (2-1) (Sugahara et al. 2021; Hashimoto et al. 2023).
@@
B14-65666 is one of the few galaxies identified as undergoing a major merger during the EoR. This system allows the investigation of the physical properties of an early major merger, including its stellar structure, as well as the different phases of the interstellar medium (ISM). Additionally, this system provides a reference for direct comparisons with simulations. In this paper, we present the first spatially resolved spectroscopy of the H emission in the system B14-65666 at â=â7.15 based on new MIRI/JWST integral field spectroscopy. The paper presents several of the physical properties of B14-65666 based on the combination of the new MIRI data with ancillary NIRCam imaging and ALMA spectroscopy. This paper is organized as follows. In Sect. 2, we present the observations and data calibrations. In Sect. 3, we outline our analysis methods of spectroscopy and photometry. In Sect. 4, we present the results of the analysis. In the subsections, we study the spectral energy distribution of the galaxy (4.1), the attenuation (4.2), the star formation rate and burstiness (4.3), the H EW and ionizing photon production efficiency (4.4), the kinematics (4.5), the stellar and dynamical mass (4.6), the mass-size relation (4.7), the starburst nature (4.8), and finally the origin and evolution of the system according to simulations (4.9). We discuss and analyze the spectral and spatial distribution of the H line and the derived properties of the galaxies, and we study the morphology and SFR of the galaxy and compare it with local galaxies. Our conclusions are presented in Sect. 5.
In this paper, we assume a flat CDM cosmology with â=â0.310, and â=â67.7âkmâs-1âMpc-1 (Planck Collaboration et al., 2020). With these parameters, 1ââ is equal to 5.275 kpc at â=â7.15 and the luminosity distance is =â72276.9 Mpc. For these cosmological parameters and redshift, the age of the Universe corresponds to 740 Myr. We use vacuum emission-line wavelengths throughout the paper.
@@ 2.1 MIRI
B14-65666 was observed with MIRI on 30 April, 2023 as a part of the MIRI European Consortium guaranteed time observations (program ID 1284, PI: L. Colina). The observations are composed of MIRI imaging (MIRIM, Bouchet et al. 2015; Dicken et al. 2024) and integral field spectroscopy using the Medium Resolution Spectrograph (MRS; Wells et al. 2015; Argyriou et al. 2023).
@@ 2.1.1 MIRI spectroscopy
The MRS observations were performed with the SHORT band for channels 1âââ4. We only used Channel 1 in this study. This channel covers the wavelengths 4.9âââ5.74âm, thus targeting the H emission line at redshift 7.15 and other fainter emission lines such as [NâII]â6550,6585â and [SâII]â6718,6733â. The total on-source exposure time of the MRS observations is 22743 seconds. The positive and negative four-point dither patterns for point sources were used. For each dither position, a total of six integrations were obtained using the SLOWR1 readout mode with 19 groups each. When the MRS targeted B14-65666, a parallel field was imaged using the MIRI F770W filter. This imaging data is utilized exclusively for the astrometric calibration of the MRS (see Sect. 2.4).
We used version 1.15.1 of the JWST calibration pipeline and context 1276 of the Calibration Reference Data System (CRDS) to process and calibrate the MRS observations of B14-65666. The standard MRS pipeline procedure (Bushouse et al. 2024) was used with additional customized steps to improve the quality of the final MRS-calibrated products and background (see Ălvarez-MĂĄrquez et al. 2023 and Ălvarez-MĂĄrquez et al. 2023 for details). Additionally, we implemented the algorithm developed by Spilker et al. (2023) to correct the striping pattern produced by the cosmic-ray showers in the spectral cube. The final Channel 1 cube has a spatial and spectral sampling of 0.13âââ0.13âââ0.8ânm (Law et al. 2023), and the resolving power () is about 3500 (Labiano et al. 2021; Jones et al. 2023).